OpenStudy (anonymous):

write an inequality to go with the following word problem. Mandy's monthly earnings consist of a fixed salary of $2,800 and an 18% commission on all her monthly sales.What must Mandy's sales be if she needs to earn an income of $6,400 per month? Please also explain how you made this into an inequality

OpenStudy (anonymous):

So she needs to make 6400 - 2800 = 3600 extra to get 6400 per month we need to find sales at 18% = 3600 or (sales)0.18 = 3600 (sales) = 3600/0.18 = 20000 so she needs to make (sales) which is 20000 to make it into an equality note that she can earn 6,400 per month or GREATER so she needs sales of \[sales \ge 20000\] or if sales is s and commision is c in percentage, i is the income she needs and w is her current wage \[s \ge \frac{ (100)(i - w) }{ c }\]